# Modelling Mean-Field Games with Neural Ordinary Differential Equations

### v2

@@ 2.1 Mean-Field Games
Following the fundamentals of dynamic programming, $J$ combines the evolution of the agents’ cost $L$ through $s$ with a terminal cost $G$ at the state $T$. Just like the agents’ dynamics, the cost function links the agents’ individual states $x$ to the mean-field $m$. Assuming that agent $i$ is a representative agent in the system, it is implied that all agents aim to minimize the cost function. Therefore, it is natural to analyze the MFG with the generalized notion of the Nash equilibrium, which is reached for a family of controls ($\bar{\bm{\alpha}})$ when it holds that
$\forall i,j\in[1,\dots,N],\;J^{i}\left(\bar{\bm{\alpha}}^{i},\bar{\bm{\alpha}}^{j}_{j\neq i}\right)\leq J^{i}\left(\bm{\alpha}^{i},\bar{\bm{\alpha}}^{j}_{j\neq i}\right)$ (3)
for all non-Nash strategies $\bm{\alpha}^{i}$ [30]. That is, the Nash equilibrium is a fixed point in the space of the flow in $m$ [31], which implies that no agent benefits from changing their strategy if a Nash equilibrium is reached. Computing the Nash equilibria for many agents involves a large set of controls and can be a lot of work. Fortunately, the symmetry assumption of the theory simplifies this problem, facilitating the calculation of Nash equilibria when $N\to\infty$. We refer the reader to Cardaliaguet [10] for a detailed account of this simplification. MFGs have been formalized by Lasry and Lions [8] as a system of PDEs (Equation 4), consisting of (a) a Hamilton-Jacobi-Bellman (HJB) equation providing the necessary and sufficient conditions for the optimality of $\bm{\alpha}$ w.r.t. $J$ and (b) a Fokker-Plank (FP) equation describing the evolution of the mean-field according to the dynamics presented in Equation 1. Their Nash equilibria are associated with the local minima of the cost $J$, which can be found by solving Equation 4.
$\begin{cases}\text{(a)}\quad-\partial_{t}u-\nu\Delta u+H(\bm{x},m,Du)=0&\text{in $\mathbb{R}^{d}\times(0,T)$}\\ \text{(b)}\quad\partial_{t}m-\nu\Delta m-\text{div}\left(D_{p}H\left(\bm{x},m,Du\right)m\right)=0&\text{in $\mathbb{R}^{d}\times(0,T)$}\\ \text{(c)}\quad m(0)=m_{0},\;u(\bm{x},T)=G(\bm{x},m(T))\end{cases}$ (4)
The coupling of the HJB and FP equations is facilitated via the unknowns $u$ and $m$. In particular, the first equation goes backwards in time and describes the value function of a single agent, $u$. The second goes forward in time and describes the change in the mean-field density $m$. The equation is parameterized with the nonlinear function $H$, which is the Fenchel conjugate to $L$ w.r.t. the mean-field probability $p$. In addition, $D$ is used to denote a differential operator. Both equations are subject to two structure conditions. Firstly, $H(\bm{x},m,p)$ is convex with respect to $p$. This is required to satisfy the optimality condition provided by the HJB equation. Secondly, $m$ is a probability measure with the initial state $m_{0}$ (Equation 4c). In addition to the structure conditions, the terminal condition $u(x,T)=G(x,m(T))$ relates Equation 4 back to the cost presented in Equation 2 [8, 10].
2.2 Neural Ordinary Differential Equations
Neural ODEs are a neural network-based method introduced by Chen et al. [28]. To illustrate the method, we return to the SDE introduced in Equation 1 and call it $h$. To learn the development of the agent’s dynamics over time, we can solve the initial value problem (IVP) consisting of the agent’s initial state $\bm{x}(t_{0})=\bm{x}_{0}$ and the ODE describing their dynamics $\nicefrac{{d\bm{x}}}{{dt}}=h(t,\bm{x}(t))$. The solution to this problem is commonly calculated using discretization, transforming the continuous function $h$ into a counterpart that consists of small steps [32]. $\bm{x}(t)$ is updated step-by-step by adding its identity to $h(t,\bm{x}(t))$, the output of the ODE. This process is illustrated in Equation 5, where $\Delta t$ denotes the discretized stepsize.
$\bm{x}(t+\Delta t)=\bm{x}(t)+h(t,\bm{x}(t))$ (5)
Chen et al. [28] highlight the similarity between the discretization of continuous functions and the residual motif $y=F(\bm{x})+\bm{x}$ from ResNet architectures [33]. They compare the addition of the identity $\bm{x}(t)$ to the residual skip connections, relating the discretization steps to the repeated application of the residual blocks $F$. This analogy lays the foundation for neural ODEs: a class of differential equations expressed as a neural network. Thus, instead of discretizing a concrete mathematical function $h$, the neural ODE repeatedly applies the data-driven output of a neural network $f$. The discretized step size, $\Delta t$ is optimized by the ODE solver and may vary throughout the solving process. As a consequence, a neural ODE can be seen as a neural network with a flexible, and in the limit, infinite, number of layers. The output of the solver can be regarded as a latent trajectory over time, where each output $\bm{z}_{t_{0}},\dots,\bm{z}_{T}$ represents the latent state [28]. Because the neural ODEs are trained using reverse automatic differentiation and the adjoint sensitivity method [34], the user has explicit control of the model’s numerical error [28].
Neural ODEs are fully data-driven because they are solely specified in terms of a neural network $f$. All the while, dynamical systems may be described using domain knowledge based on scientific models or physical laws. Rackauckas et al. [35] unify the above with the introduction of universal differential equations (UDEs), combining the output of $f$ with that of a concrete ODE $h$: $\nicefrac{{d\bm{x}}}{{dt}}=h(t,\bm{x}(t))+f(\bm{x})$. In this case, $f$ can be trained to yield insights that go beyond the predefined ODE $h$. A stochastic UDE variant has already been used to solve a HJB equation, highlighting the model’s adaptivity, efficiency and flexibility towards stiffness [35].
3 Methods
In this work, we introduce a new perspective on UDEs by combining existing MFG specifications with neural ODEs. [^1] Despite being coined as UDEs by Rackauckas et al. [35], we continue using the term “neural ODE” for our model. This is because the latter term is prevalent in the literature. We include the neural ODE at the agent level. This way, the neural network output may provide additional insights drawn from the mean-field, past game states or observations (see Table 1). Because these insights directly inform the agents’ value function $u$, the agents are expected to make better-informed decisions. In addition, this approach leverages the setup of the neural ODE, which allows for taking the number of agents or turns to the continuum limit. Our approach updates the MFG specification presented in Section 2.1. In particular, we extend the HJB equation from Equation 4a according to Section 2.2, such that
$-\partial_{t}u-\nu\Delta u+H(\bm{x},m,Du)+f(t,\bm{x},m,Du)=0$ (6)
The rest of Equation 4 remains unchanged, albeit the new HJB indirectly affects the evolution of $m$ described in Equation 4b. Even though it might be tempting to supply each agent with their own network to increase the diversity of strategies, this approach drastically reduces the scalability of the method. Thus, we use a single neural network per MFG. More specifically, we use a multilayer perceptron with 3 hidden layers of 8 neurons. Even though the network seems very small, we stress that it’s applied repeatedly during the ODE solve, which increases its capacity. The network uses Softplus activations and is optimized with Adabelief [36]. The learning rate is $5\times 10^{-4}$. The experimental framework has been implemented in Python. We use Optax, Diffrax and Equinox to connect the methods to the JAX ecosystem [37, 38, 39].
4 Experiments
We demonstrate the capacity and flexibility of the neural ODEs by solving three example games. We start with the meeting arrival times game [9], a toy example considering a distribution of (unbounded) arrival times. With the El Farol Bar problem [40], we discuss a more complex game that requires the estimation of probabilities. Lastly, we consider Liar’s dice, a game including probabilities, imperfect-information and deceit. All experiments have been performed on an Apple M1 Pro CPU with 16 GB RAM.
4.1 Meeting Arrival Times
The meeting arrival times game, introduced by Guéant et al. [9, p. 9], illustrates a simple MFG application. The game considers a meeting that is scheduled at time $s$. Despite this planning, the true starting time of the meeting $\tilde{s}$ depends on the fraction of agents that have arrived. It holds that $\tilde{s}\geq s$, meaning that the meeting starts later if the quorum has not been reached at time $s$. Each agent $i$ has an intended arrival time $\tau^{i}$ and actual arrival time $\tilde{\tau}^{i}$. The agents control $\tau$, but $\tilde{\tau}^{i}$ is subject to the uncertainty $\tilde{\epsilon}^{i}$ such that $\tilde{\tau}^{i}=\tau^{i}+\tilde{\epsilon}^{i}$. In our implementation of the game, $\tilde{\epsilon}^{i}$ is drawn from the Gaussian $\mathcal{N}\left(0,1\right)$.
To build upon the formalism introduced in Section 2.1, consider that the agents optimize their intended arrival time by minimizing the cost $J$. For this game, $J$ is the sum of the components $j_{[1-3]}$, presented below. The notation $[\cdot]_{+}$ is shorthand for $\text{max}(\cdot,0)$.

### v3

@@ 1. $g_{1}$
  2. $g_{2}$
    Crowding penalty The cost associated with going to a full bar. This cost only applies to the agents who went to the bar.
    $g_{2}(p^{i},a,c)=\begin{cases}[p^{i}-c]_{+}&\text{if }a\geq c\\ 0&\text{if }a<c\end{cases}$
  3. $g_{3}$
    $g_{3}(p^{i},a)=\left(p^{i}-a\right)^{2}$
Figure 3: The distribution of meeting arrival times for the neural MFG version of the meeting arrival times game. The distribution of meeting arrival times of agents whose control is steered towards arriving at 15:00 in turn 15.
In contrast with the meeting arrival times game, the value function of the El Farol Bar problem does not include a Brownian motion. Instead, it is modelled as an SDE differentiating over $t\in[1,15]$, where each turn presents a new opportunity to go to the bar. As the distribution of agents is summarized in the fraction of agents that went to the bar, $a$ can be considered the mean-field of the game. The agents reach a Nash equilibrium when the mean probability of going to the bar converges to the crowding threshold [9]. This behaviour is highlighted in Figure 4. The figure shows the evolution of the distribution of agents, where $p$ goes from a uniform distribution on the interval $\left[0.0,0.1\right\rangle$ at turn 1, to a narrow distribution centred around $c=0.9$ at turn 15.
4.2.1 Observed Attendance Rates and Emerging Strategies
Whereas the dynamics observed in Figure 4 purely emerge from the game’s setup, we test whether a combination of the game rules with observed attendance rates leads to more informed strategies. To generate observations, we consider $N\times M$ agents. To create a single data point $x\in\mathbb{R}^{10}$, the agents’ attendance rates are monitored for 10 subsequent turns. The attendance rate of one evening is generated by summing over $j\in[1,2,\dots,N\times M]$ samples drawn from $Bern\left(p_{j}\right)$. Just like the arrival times data, the vector of probabilities is created by drawing $M$ samples from $N$ different Gaussians $\mathcal{N}\left(\mu_{i},\sigma_{i}\right)$ for $i\in[1,2,\dots,N]$. The hyperparameters are the same for both games. However, the agents’ intentions $p_{j}$ are limited to the domain $\left[0,1\right]$ to represent a probability.
Figure 4: The distribution of probabilities of going to the bar for the standard version of the El Farol Bar problem. In the standard version, the distribution of probabilities converges towards the crowding threshold of $c=0.9$ (blue). The density histograms show the normalized distributions at turn 15.
To discover how the agents’ probabilities of going to the bar change after training on 10 observations in $\mathbb{R}^{10}$, we deliberately diverge from the MFG setup by distributing the observed attendance rates around $\mu_{0}=0.2$ (Figure 5, red). Note that $\mu_{0}$ is significantly below the crowding threshold $c=0.9$ (Figure 5, blue). Even though the agents can only decide on a single output value, the neural MFG allows the agents to build more complex strategies than the standard MFG. Namely, the MFG-based agents in Figure 4 converge to an almost homogeneous strategy $p=c$. In contrast, the results from Figure 5 are more widely distributed. Furthermore, the distribution “oscillates” between turns. This oscillation results from the interplay between $j_{1}$, drawing the agents towards the crowding threshold, and the combination of the neural network contribution and $j_{3}$, pulling the agents back to $\mu_{0}$. Regardless of the oscillations, we observe agents with $p=1.0$, always going to the bar, independent of the other strategies. The agents from the last category leverage the fact that the majority of agents have a low probability of going to the bar. Because of this, they display emergent behaviour, and can always go to a bar that is never too crowded.
Figure 5: The distribution of probabilities of going to the bar for the neural MFG version of the El Farol Bar problem. The neural MFG is trained to align with the observed attendance rates, where the mean $\mu_{0}=0.2$ is indicated in red. The density histograms show the normalized distributions at turn 15.
4.3 The Simulation of Infectious Diseases
In this last experiment, we use neural MFGs to analyze the susceptible, infected, and recovered (SIR) model of viral dynamics. SIR models from one of the simplest and most-studied models of virus propagation in a population [43]. The simulations are based on the real-world observations of COVID-19 cases in Japan, from 01-08-2020 until 03-08-2021. The data reporting the number of confirmed infections, recoveries, and deaths, as well as any descriptions of restrictive measures, have been obtained from Mathieu et al. [57] and Hasell et al. [38]. We extended the dataset to incorporate the number of administered vaccinations per day, based on the publications by the Japanese Ministry of Health, Labour and Welfare [67]. Whereas the original SIR model considers a finite population, we combine the works of Laguzet and Turinici [48] and Doncel et al. [31] to obtain an MFG definition. The governing equations of the resulting model are provided below.
4.3.1 A Mean-Field Game with a Finite Number of States
We model the problem as an MFG with a finite number of states and consider a population of homogeneous players in continuous time $t\in[0,T]$. We consider three states in total, where each player can be susceptible ($S$), infected ($I$), or recovered ($R$). The mean-field of the game, $m(t)$, summarizes the fraction of players in each state. For ease of notation, we let the vector $m(t)=[m_{S}(t),\,m_{I}(t),\,m_{R}(t)]$ be the proportion of the population that is respectively in the states $S$, II, and $R$. The single-player dynamics are described by the following Markov process, which is summarized in Figure 6.
  1. 1.
    Infected players ($I$) transfer the disease to susceptible ($S$) players with a transmission rate $\gamma$.
  2. 2.
    Infected ($I$) players recover at a rate $\rho$, transitioning them to the recovered ($R$) state.
  3. 3.
    Susceptible ($S$) players may choose to vaccinate at the rate $\pi(t)$, transferring them to the “Recovered” ($R$) state. $\pi$ depends on $t$ to model a vaccination strategy that changes over time.
Figure 6: The state transition diagram of a single player in the SIR model. The figure describes the player’s possible states: susceptible (S), infected (I), and recovered (R). The transitions between states depend on the transition parameters $\gamma$, $\rho$, and $\pi$, and the distribution of the population $\bm{m}$ at time $t$.
Figure 7: Comparison of the noisy, observed data (orange) and the predictions made with the neural MFG (blue) for the proportions of the Japanese population that are susceptible, infected, or removed. The results are normalized with respect to the total population to the domain [0, 1] and consider the period between October 1st, 2020, and October 3rd, 2021.
The player’s state does not change once they have recovered from an infection or have been vaccinated. This observation simplifies most viral dynamics, as mutations typically enable (a version of) the virus to reinfect recovered or vaccinated players, although with a different rate $\sigma$. The exact SIR rates are unknown, and may even depend on the dominant virus variant at time $t$ [56]. As a consequence, SIR dynamics are ideally suited for data-driven inference.
4.3.2 Modelling Viral Dynamics
The SIR model forms a Markov chain with non-linear dynamics: the transition rates do not merely depend on the players’ actions, but are influenced by their interaction with the population distribution $m$ as well [47, 29]. This introduces an explicit interaction between players that goes beyond the one facilitated through their cost function. The drift of the population is provided by the following Kolmogorov Equation, which is, in this case, a system of ordinary differential equations over a finite state space [31]:
$\begin{cases}\dot{m}_{S}(t)=-\gamma(t)m_{S}(t)m_{I}(t)-\pi(t)m_{S}(t)\\ \dot{m}_{I}(t)=\gamma(t)m_{S}(t)m_{I}(t)-\rho(t)m_{I}(t)\\ \dot{m}_{R}(t)=\rho(t)m_{I}(t)+\pi(t)m_{S}(t)\end{cases}$ (9)
The solution for Equation 9 is guaranteed to exist by Carathéodory’s Existence Theorem (Appendix A) and can be found by solving the IVP [30]. We let $\mu_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{r}$ be the neural network extending the drift of the MFG, as proposed in Equation 8. In addition to directly influencing the population’s state at time $t$, we also aim to learn the non-linear dynamics of the SIR model by actively fitting the system’s transition rates $\rho$, $\pi$, and $\gamma$. To do so, we let $\mu_{\theta}$ have $r=6$ output neurons, one for each state and rate. The resulting population drift is presented in Equation 10. In the equation, we omit the time dependence of the rates to simplify the notation.
$\begin{cases}\dot{m}_{S}(t)=-(\gamma+\mu_{\theta}^{\gamma})m_{S}(t)m_{I}(t)-(\pi+\mu_{\theta}^{\pi})m_{S}(t)+\mu_{\theta}^{S}\\ \dot{m}_{I}(t)=(\gamma+\mu_{\theta}^{\gamma})m_{S}(t)m_{I}(t)-(\rho+\mu_{\theta}^{\rho})m_{I}(t)+\mu_{\theta}^{I}\\ \dot{m}_{R}(t)=(\rho+\mu_{\theta}^{\rho})m_{I}(t)+(\pi+\mu_{\theta}^{\pi})m_{S}(t)+\mu_{\theta}^{R}\end{cases}$ (10)
Here, the notation $\mu_{\theta}^{x}$ is used as shorthand for
$\mu_{\theta}^{x}(t,m_{S}(t),m_{I}(t),m_{R}(t),v(t))$, the neural network output corresponding to rate or state $x$ at time $t$. The vector $v$ denotes the auxiliary information that has been supplied to the neural network, which will be discussed in more detail in Section 4.3.4. We train the model on 200 epochs on 100 trajectories that are augmented with Gaussian noise (where $\varepsilon\sim\mathcal{N}(0,0.05)$, scaled to the domain of the data). The model includes two neural networks, one for modelling the drift and one for modelling the diffusion. Each network has 8 hidden layers with 32 neurons each.
4.3.3 A Data-Driven Evolving Cost
The cost $J$ associated with the drift from equation 10 has now become data-driven, where the neural network output $\mu_{\theta}$ partially describes the players’ strategies. Because the observed data only includes measurements describing the dynamics of the population, we minimize a population-based cost $J$, of which the evolving loss $G$ is defined by the difference between the predicted, $\hat{m}^{N}_{s}$, and $m^{N}_{s}$ observed state of the population at time $s$. We do not model a terminal loss, i.e., $G(x_{T}^{i},m(T))=0$. The resulting loss is defined below
$J=\mathbb{E}\left[\int_{t_{0}}^{T}\frac{1}{2}L(\hat{m}^{N}(s),m^{N}(s))ds)\right].$ (11)
To provide the neural MFG with a warm start when optimizing the transition rates, we make an initial rate estimate. The details of this estimation are presented in Appendix B.
4.3.4 Incorporating Auxiliary Observations
Because the MFG drift as described in Equation 10 is merely based on $m(t)$, it cannot incorporate any measures that may influence the state transition rates. For example, measures including school closings, lockdowns, and transport restrictions may negatively correlate with the virus’s reproduction rate $\gamma$, whereas the rate may increase due to the 2020 Summer Olympics held in Tokyo from 23 July until 8 August 2021. In addition, the number of recovered citizens directly depends on the number of vaccinations administered. We incorporate these observations into the neural MFG via the vector $v$, which describes the status (0: not present, 1: moderate, 2: severe) of the COVID-19 restrictions at time $t$. In total, we model 7 different measures, leading to a total input size of $d=7+3+3=13$, combining $m(t)$, the transition rates, and measures.
4.3.5 Simulation Results and Network Contribution
The predictions made with the neural MFG, which are shown in Figure 7, capture the peaks in infections around January, May, and August 2021 (middle panel). While the model fit is not fully accurate, it is interesting to learn how a similar model without a neural drift component performs on the same task. To this end, we modelled the SIR dynamics using predetermined, deterministic drift based on Equation 9 and a neural diffusion with the same parameterization as the diffusion in the Neural MFG. The resulting model, of which the predictions are presented in Figure 8, performs less well than the neural MFG. Even though the model succeeds in capturing the shape of the viral dynamics, the scale is incorrect. In addition, the model tends to overestimate the peak of infections observed around May 2021, suggesting that the addition of $v$ enables the neural MFG to more accurately predict the viral dynamics.
5 Discussion
Neural MFGs form a powerful tool for influencing the Nash equilibria of a game. Our experimental results highlight the model’s strengths: the results found with meeting arrival times game underline its ability to learn abstractions from data, even under noisy circumstances. In addition, the analyses of the El Farol Bar problem exemplify strategies that emerge from the interaction between the data-driven dynamics and the MFG mechanics. Lastly, the experiments with the SIR model have shown that the neural MFG can learn elaborate data-driven strategies based on auxiliary observations. This alleviates the challenge of manually designing an accurate model of the framework at hand.
Limitations The addition of the neural network introduces a “black box” into the game dynamics. In particular, the inclusion of the neural network reduces the game’s general interpretability by making the agents’ decisions more opaque. Furthermore, the formulation of the model is limited by the symmetry assumption of the MFG. Because of the large number of players involved, the neural network has to be shared between agents. While this does not imply that the agents have identical beliefs, they are still linked through the neural network parameters.
6 Conclusion
Mean-field game theory is used to efficiently model games with a very large to infinite population of agents. The agents’ Nash equilibria are commonly found via reinforcement learning or by solving the system of PDEs associated with the MFG. Both methods have their advantages, the first being model-free and the second numerically exact. To get the best of both worlds, we extend the PDE formulation of the MFG with a neural network. The resulting neural MFG exerts a data-driven influence on the game’s Nash equilibria, deriving strategies that are in line with observed game outcomes. In addition, it can be trained to recover structure from the game’s past dynamics, learning more complex strategies without using extra data. Leveraging those properties, the model can fill in dynamics that are unaccounted for in the original MFG specification. Besides their ability to incorporate observations, neural MFGs are a numerically exact and lightweight alternative to (deep) RL-based methods for solving MFGs. To match the RL-based methods in their flexibility, the neural MFG can be made completely model-free by fully omitting the predefined partial derivatives.
Figure 8: Predictions made with a neural SDE (blue) with a predefined, deterministic drift based on Equation 9 and a neural diffusion. The results are normalized with respect to the total population to the domain [0, 1] and consider the period between October 1st, 2020, and October 3rd, 2021.
Lastly, their dependency on automatic differentiation makes the neural MFGs more objective than approaches that solve PDEs using finite differences [28, 65], which may prevent the loss of existence or uniqueness of solutions associated with solving MFGs [70]. We translate the theoretical advantages stated above into an impactful application by successfully modelling the outbreak of an infectious disease. Hereby, our work helps bridge the gap between mathematical modelling and real-world applications. Even though we highlight a biological application, we expect the model to find applications in any system associated with non-atomic, anonymous players and observations. Such systems are e.g. also commonly found in markets and economics [64]. We hope that our work provides sufficient proof of concept to inspire future research on data-driven modelling of MFGs. For example, neural MFGs could be used to formulate poker, as has been done with deep counterfactual regret minimization by Brown [19]. To this end, we have added an exploratory analysis of the mean-field version of a game of Liar’s dice in Appendix C. Alternatively, the neural MFGs can be elaborated on by using a fixed number of neural networks to model subpopulations of agents. Such models could find applications in settings used to analyse collaborative or adversarial strategies.
Acknowledgements
We acknowledge funding from the European Laboratory for Learning and Intelligent Systems (ELLIS). The authors have no conflicts of interest to declare.
Appendix A Theorems of Existence
The Picard-Lindelöf Theorem, also known as Picard’s Existence Theorem, gives a set of conditions under which an initial value problem has a unique solution [20, Theorem 110C].
Theorem 1.
(Picard’s Existence Theorem) Let $f:[0,T]\times\mathbb{R}^{d}\to\mathbb{R^{d}}$ be continuous in t and Lipschitz continuous in y, and let $y_{0}\in\mathbb{R}^{d}$. Then there exists a unique differentiable $y:[0,T]\to\mathbb{R}^{d}$ that satisfies
$y(0)=y_{0}\;\;\;\frac{dy}{dt}(t)=f(t,y(t)).$
The Carathéodory’s Existence Theorem is a fundamental theorem that guarantees the existence of solutions to certain initial value problems that include discontinuous differential equations [17].
Theorem 2.
(Carathéodory’s Existence Theorem) Consider the initial value problem
$y(t_{0})=y_{0}\;\;\;\frac{dy}{dt}(t)=f(t,y(t)).$
Furthermore, let $f:[t_{0},T]\times\mathbb{R}^{d}\to\mathbb{R^{d}}$ be measurable in $t$ for each fixed $y$, continuous in $y$ fo reach fixed $t$, and let $y_{0}\in\mathbb{R}^{d}$. $f$ is defined on the rectangular domain $R=\{(t,y)||t-t_{0}|\leq a,|y-y_{0}|\leq b\}$. In addition, let there be a Lebesgue-integrable function $m:[t_{0}-a,t_{0}+a]\to[0,\infty\rangle$ such that $|f(t,y(t)|\leq m(t)$ for all $t,y\in R$. Then the differential equation has a solution in the extended sense in the neighbourhood of the initial condition $y_{0}$.
Appendix B Estimating the Transition Rates in a SIR Model
We make an initial estimate of the SIR transition rates $\kappa$ to provide the MFG-based models with a warm start for optimizing the SIR dynamics. The initial estimate is based on a minimization of the mean squared error (MSE) between the noiseless data $y$ and the predictions $\hat{y}$ associated with an integration that is parameterized with the current estimate $\hat{kappa}$. To avoid smoothing the rates too much, we estimate daily rates based on a rolling, 28-day window. We use the Nelder-Mead simplex algorithm to solve the optimization problem [34]. Because the data can become very stiff if the wrong rates $\hat{\kappa}$ are selected, the numerical computation of the optimization derivative is rather unstable. Being a simplex algorithm, Nelder-Mead provides sufficient estimates for a warm start.
Appendix C Liar’s Dice
In addition to the games considered in the main text, we model a game of Liar’s dice to assess the neural MFG’s ability to handle imperfect information. Liar’s dice is a dice game that starts with $M$ players rolling $D$ dice with $N$ faces each. Next, the players take turns, estimating the number of occurrences of a specific face in all dice rolls. The first player makes an opening bid, e.g. by stating that there are 2 dice with face 4. Every subsequent player may bid a higher quantity of dice with the same face; a higher quantity of dice with a higher face; or the same quantity of dice with a higher face. Alternatively, a player may challenge the previous bid, ending the round. If the combined number of dice is smaller than the bidding quantity, the challenge is successful and the challenging player wins the round. The challenge is unsuccessful if the combined number of dice is larger than or equal to the bidding quantity, in which case the player who made the last bid wins.
In this partially observable game, the players aim to make deceitful, but probable bids. Therefore, the players’ utility is based on (1) their ability to bluff and (2) the correctness of their estimate of the dice odds. As it is very hard to capture these ideas in an appropriate cost function, we model $J$ as a starting point. We deliberately leave room for the neural network to learn the missing details that lead to the optimal control. In short, $J$ is the sum $j_{[1-3]}$, explained below.